xx HydroAnalysis

HydroxAnalysis

Frank Lxffler

May 17, 2024

Abstract

Provides basic hydro analysis routines and quantities.

1 Purpose

2 Parameters

11hydroxanalysisxaveragexmultiplexmaximaxlocations 10Scope: privatex01  BOOLEAN

11Description: when finding more than one global maximum location, average position and use result

11                                         10  x01  Default: false

11

11hydroxanalysisxcompxcorexrhoxcentroid  10Scope: privatex01  BOOLEAN

11Description: compute location of the centroid of rho*x in region rxcore around densest point

11                                         10  x01  Default: no

11

11hydroxanalysisxcompxrhoxmax  10Scope: privatex01  BOOLEAN

11Description: Look for the value and location of the maximum of rho

11                                         10  x01  Default: false

11

11hydroxanalysisxcompxrhoxmaxxevery  10Scope: privatex01  INT

11Description: How often to look for the value and location of the maximum of rho

11 Range 10 x01 Default: 1
110:0  10Never
111:*  10every so often

11

11hydroxanalysisxcompxrhoxmaxxoriginxdistance 10Scope: privatex01  BOOLEAN

11Description: Look for the proper distance between the maximum of the density and the origin (along a straight coordinate line)

11                                         10  x01  Default: false

11

11hydroxanalysisxcompxvolxweightedxcenterxofxmass 10Scope: privatex01  BOOLEAN

11Description: Look for the location of the volume-weighted center of mass

11                                         10  x01  Default: false

11

11hydroxanalysisxcorexrhoxrelxmin  10Scope: privatex01  REAL

11Description: only include points where rho>rhoxrelxmin*rhoxmax when computing centroid of rho*x around densest point

11 Range 10 x01 Default: 0.01
110:*  10any positive value. For best results should be such that the region selected is smaller than rxcore. 1/10 of rhoxmax should be fine.

11

11hydroxanalysisxinterpolatorxcoordinates  10Scope: privatex01  STRING

11Description: Coordinate system

11 Range 10 x01 Default: cart3d
11.*  10must be a registered coordinate system

11

11hydroxanalysisxinterpolatorxname  10Scope: privatex01  STRING

11Description: Name of the interpolator

11 Range 10 x01 Default: uniform cartesian
11see [1] below  10from AEILocalInterp
11see [1] below  10from AEILocalInterp
11see [1] below  10from AEILocalInterp
11uniform cartesian  10from LocalInterp
11.*  10must be a registered interpolator

11

Lagrangexpolynomialxinterpolationx(tensorxproduct)

[1]

Lagrangexpolynomialxinterpolationx(maximumxdegree)

[1]

Hermitexpolynomialxinterpolation

11hydroxanalysisxinterpolatorxoptions  10Scope: privatex01  STRING

11Description: Options for the interpolator

11 Range 10 x01 Default: order=2
11.*  10must be a valid option specification

11

11hydroxanalysisxrxcore  10Scope: privatex01  REAL

11Description: size of region around densest point in which to compute the centroid of rho*x

11 Range 10 x01 Default: 4.0
110:*  10any positive radius. suggested is 1/2 of stellar radius

11

11hydroxanalysisxrhoxmaxxlocxonlyxpositivexx 10Scope: privatex01  BOOLEAN

11Description: Restrict location search for density maximum to positive values of x

11                                         10  x01  Default: false

11

11hydroxanalysisxrhoxmaxxlocxusexrotatingsymmetry180 10Scope: privatex01  BOOLEAN

11Description: Map found maxima into positive x half-plane assuming pi-symmetry

11                                         10  x01  Default: false

11

11hydroxanalysisxrhoxmaxxoriginxdistancexnpoints10Scope: privatex01  INT

11Description: Number of points along the straight line for measuring proper distance

11 Range 10 x01 Default: 100
111:*  10Any positive number

11

11verbosityxlevel  10Scope: privatex01  INT

11Description: how much information to ouptut to the logs

11 Range 10 x01 Default: 1
11                                         10Output nothing
111  10warn when finding multiple maxima
112  10also output location of maxima

11

11restmassxcomputexmasses  10Scope: restrictedx01  BOOLEAN

11Description: Should we compute the masses?

11                                         10  x01  Default: no

11

11restmassxmassesxnr  10Scope: restrictedx01  INT

11Description: number of radii within which to compute the rest mass

11 Range 10 x01 Default: (none)
110:100  10Positive and hard-limited

11

11restmassxrefxradiusxmass  10Scope: restrictedx01  REAL

11Description: Radii within which mass will be computed

11 Range 10 x01 Default: 10.0
110.0:  10Positive

11restmassxrhoxmin  10Scope: restrictedx01  REAL

11Description: Points with rest-mass density below this value are excluded (it could be set equal to the atmosphere value)

11  Range  10  x01  Default: 1.e-9

110:  10

11

11timelevels  10Scope: shared from HYDROBASEx01INT

11

3 Interfaces

General Implements: hydroxanalysis

Inherits: grid hydrobase admbase

Grid Variables

3.0.1 PRIVATE GROUPS

11  x Group Names x  10  x Variable Names x  x01 Details x11x
11  gridxspacingxproduct  10  gridxspacingxproduct  x01  compact110
11                           10  x01 description11product of cctkxdeltaxspace
11 10 x x01 description11to be computed in local mode and later used in global mode
11                           10  x01 dimensions110
11                           10  x01 distribution11CONSTANT
11                           10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                           10  x01 timelevels111
11                           10  x01variable type11REAL
11 hydroxanalysisxmasses 10 HydroxAnalysisxmasses x01 compact110
11                           10  x01 description11Baryonic masses at different radii
11                           10  x01 dimensions110
11                           10  x01 distribution11CONSTANT
11                           10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                           10  x01 timelevels111
11                           10  x01vararrayxsize11restmassxmassesxnr
11                           10  x01variable type11REAL
11hydroxanalysisxmassesxfractions10HydroxAnalysisxmassesxfractionsx01  compact110
11                           10  x01 description11Fractional Baryonic masses at different radii
11                           10  x01 dimensions110
11                           10  x01 distribution11CONSTANT
11                           10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                           10  x01 timelevels111
11                           10  x01vararrayxsize11restmassxmassesxnr
11                           10  x01variable type11REAL
11 hydroxanalysisxtotalxrestxmass 10  x01  compact110
11                           10  totalxrestxmass  x01 description11Total Baryonic mass
11                           10  x01 dimensions110
11                           10  x01 distribution11CONSTANT
11                           10  x01 group type11SCALAR
11 10 x01 tags11checkpoint=”no”
11                           10  x01 timelevels111
11                           10  x01variable type11REAL
11 hydroxanalysisxmassesxtemps 10 HydroxAnalysisxmassesxtemps x01  compact110
11                           10  x01 description11Temporaries for the mass calculation
11                           10  x01 dimensions113
11                           10  x01 distribution11DEFAULT
11                           10  x01 group type11GF
11 10 x01 tags11checkpoint=”no”
11                           10  x01 timelevels113
11                           10  x01vararrayxsize11restmassxmassesxnr+1
11                           10  x01variable type11REAL

11

3.0.2 PUBLIC GROUPS

11  x Group Names x  10  x Variable Names x  x01 Details x11x
11  hydroxanalysisxrhoxmax  10  HydroxAnalysisxrhoxmax  x01  compact110
11                                        10  x01 description11value of the maximum of rho
11                                        10  x01 dimensions110
11                                        10  x01 distribution11CONSTANT
11                                        10  x01 group type11SCALAR
11                                        10  x01 timelevels111
11                                        10  x01variable type11REAL
11 hydroxanalysisxrhoxsum 10 HydroxAnalysisxrhoxsum x01 compact110
11                                        10  x01 description11value of the sum of rho
11                                        10  x01 dimensions110
11                                        10  x01 distribution11CONSTANT
11                                        10  x01 group type11SCALAR
11                                        10  x01 timelevels111
11                                        10  x01variable type11REAL
11 hydroxanalysisxrhoxmaxxloc 10 HydroxAnalysisxrhoxmaxxloc x01 compact110
11                                        10  x01 description11coordinate location of the maximum of rho
11                                        10  x01 dimensions110
11                                        10  x01 distribution11CONSTANT
11                                        10  x01 group type11SCALAR
11                                        10  x01 timelevels111
11                                        10  x01vararrayxsize113
11                                        10  x01variable type11REAL
11hydroxanalysisxrhoxcenterxvolumexweightedxgf10HydroxAnalysisxrhoxcenterxvolumexweightedxgfx01  compact110
11                                        10  x01 description11temporary GF to obtain the coordinate location of the volume weighted center of mass via a reduction
11                                        10  x01 dimensions113
11                                        10  x01 distribution11DEFAULT
11                                        10  x01 group type11GF
11 10 x01 tags11checkpoint=”no” tensortypealias=”u”
11                                        10  x01 timelevels113
11                                        10  x01vararrayxsize113
11                                        10  x01variable type11REAL
11 hydroxanalysisxrhoxcenterxvolumexweighted 10 HydroxAnalysisxrhoxcenterxvolumexweighted x01  compact110
11                                        10  x01 description11coordinate location of the volume weightes center of mass
11                                        10  x01 dimensions110
11                                        10  x01 distribution11CONSTANT
11                                        10  x01 group type11SCALAR
11                                        10  x01 timelevels111
11                                        10  x01vararrayxsize113
11                                        10  x01variable type11REAL
11 hydroxanalysisxrhoxmaxxoriginxdistance 10 HydroxAnalysisxrhoxmaxxoriginxdistance x01  compact110
11                                        10  x01 description11proper distance between the maximum of the density and the origin (along a straight coordinate line)
11                                        10  x01 dimensions110
11                                        10  x01 distribution11CONSTANT
11                                        10  x01 group type11SCALAR
11                                        10  x01 timelevels111
11                                        10  x01variable type11REAL

11  x Group Names x  10  x Variable Names x  x01 Details x11x
11 hydroxanalysisxcorexrhoxsum 10 HydroxAnalysisxcorexrhoxsum x01  compact110
11                               10  x01 description11value of the sum of rho in the core region
11                               10  x01 dimensions110
11                               10  x01 distribution11CONSTANT
11                               10  x01 group type11SCALAR
11                               10  x01 timelevels111
11                               10  x01variable type11REAL
11hydroxanalysisxcorexrhoxcentroidxgf10HydroxAnalysisxcorexrhoxcentroidxgfx01  compact110
11                               10  x01 description11temporary GF to obtain the coordinate location of the centroid location of rho*x in the core region via a reduction
11                               10  x01 dimensions113
11                               10  x01 distribution11DEFAULT
11                               10  x01 group type11GF
11 10 x01 tags11checkpoint=”no” tensortypealias=”4u”
11                               10  x01 timelevels113
11                               10  x01vararrayxsize114
11                               10  x01variable type11REAL
11 hydroxanalysisxcorexrhoxcentroid 10 HydroxAnalysisxcorexrhoxcentroid x01  compact110
11                               10  x01 description11coordinate location of the centroid of rho*x in the core region
11                               10  x01 dimensions110
11                               10  x01 distribution11CONSTANT
11                               10  x01 group type11SCALAR
11                               10  x01 timelevels111
11                               10  x01vararrayxsize113
11                               10  x01variable type11REAL

11

4 Schedule This section lists all the variables which are assigned storage by thorn EinsteinAnalysis/HydroxAnalysis. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11x11Conditional:
11x 11HydroxAnalysisxrhoxmax
11x 11HydroxAnalysisxrhoxmaxxloc
11x 11HydroxAnalysisxtotalxrestxmass
11x 11HydroxAnalysisxmassesxfractions
11x 11gridxspacingxproduct
11x 11HydroxAnalysisxrhoxsum
11x 11HydroxAnalysisxrhoxcenterxvolumexweighted
11x 11HydroxAnalysisxrhoxcenterxvolumexweightedxgf[3]
11x 11HydroxAnalysisxrhoxmaxxoriginxdistance
11x 11HydroxAnalysisxcorexrhoxsum
11x 11HydroxAnalysisxcorexrhoxcentroid
11x 11HydroxAnalysisxmassesxtemps[3]
11x 11HydroxAnalysisxmasses

11x 11x

Scheduled Functions

CCTKxBASEGRID (conditional) x hydroxanalysisxinit x initialize variables x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxPOSTSTEP (conditional) x hydroxanalysis x group for hydroxanalysis routines x

11x11Type:11group

11

HydroxAnalysisxCompCoreRhoCentroid (conditional) x hydroxanalysisxcompcorerhocentroidxgetstorage x get temporary storage for duration of reduction x

11x11Before: 11hydroxanalysisxcompcorerhocentroidxpreparereduction
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

HydroxAnalysisxCompCoreRhoCentroid (conditional) x hydroxanalysisxcompcorerhocentroidxpreparereduction x prepare core center of mass densities for reduction x

11x11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Type: 11function

11

HydroxAnalysisxCompCoreRhoCentroid (conditional) x hydroxanalysisxcompcorerhocentroidxreduction x compute the global core center of mass reduction results x

11x11After: 11hydroxanalysisxcompcorerhocentroidxpreparereduction
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

HydroxAnalysisxCompCoreRhoCentroid (conditional) x hydroxanalysisxcompcorerhocentroidxfreestorage x free temporary storage x

11x11After: 11hydroxanalysisxcompcorerhocentroidxreduction
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxPOSTSTEP (conditional) x hydroxanalysisxmasses x computation of rest masses contained within given radii x

11x11Type:11group

11

HydroxAnalysisxMasses (conditional) x hydroxanalysisxmassesxlocal x rest masses within given radii, local calculations x

11x11Language:11fortran
11 x 11Type: 11function

11

HydroxAnalysisxMasses (conditional) x hydroxanalysisxmassesxglobal x rest masses within given radii, reduction operation x

11x11After: 11hydroxanalysisxmassesxlocal
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

11

HydroxAnalysis (conditional) x hydroxanalysisxpreparereduction x compute the local reduction results x

11x11Language:11c
11 x 11Options: 11global-early
11 x 11x  11loop-local
11 x 11Type: 11function

11

HydroxAnalysis (conditional) x hydroxanalysisxreduction x compute the global reduction results

11x11After: 11hydroxanalysisxpreparereduction
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

HydroxAnalysis (conditional) x hydroxanalysisxlocationsearch x look for the location of the maximum density x

11x11After:11hydroxanalysisxreduction
11 x 11Type:11group

11

HydroxAnalysisxLocationSearch (conditional) x hydroxanalysisxlocationsearchxsetup x prepare data structures for search x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

HydroxAnalysisxLocationSearch (conditional) x hydroxanalysisxlocationsearchxsearch x search for the location of the maximum density x

11x11After: 11hydroxanalysisxlocationsearchxsetup
11 x 11Language:11c
11 x 11Options: 11global
11 x 11x  11loop-local
11 x 11Type: 11function

11

HydroxAnalysisxLocationSearch (conditional) x hydroxanalysisxlocationsearchxcombine x communicate and verify the location of the maximum density x

11x11After: 11hydroxanalysisxlocationsearchxsearch
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

HydroxAnalysis (conditional) x hydroxanalysisxfindseparation x compute the proper distance between the maximum of the density and the origin (along a straight coordinate line) x

11x11After: 11hydroxanalysisxlocationsearch
11 x 11Language:11fortran
11 x 11Options: 11global
11 x 11Type: 11function

HydroxAnalysis (conditional) x hydroxanalysisxcompcorerhocentroid x compute center of mass of core region x

11x11After:11hydroxanalysisxlocationsearch
11 x 11Type:11group